# Towards Combinatorial Interpretability of Neural Computation

### v1

@@ 1 Introduction
An exciting recent body of research in the interpretability of neural networks has been focused on viewing their computation as collections of independent features whose directions in high dimensional activation space can be analyzed. This has led, using novel approaches such as the use of sparse autoencoders [31], to the discovery of polysemanticity, the superposition hypothesis [12], and a body of empirical evidence of the correlation between features and directionality of neuronal activation space vectors [2, 3, 5, 7, 10, 11, 14, 31, 32]. Those techniques (surveyed in detail in the related work section) have provided ingenious ways of capturing the representation of features and their interdependencies, but so far have not elucidated the actual underlying computational mechanisms of gradient-descent-trained neural networks.
@@ 2 Background and Related Work
Recent advances in the use of autoencoders for network interpretability have led to enhanced methods for isolating and manipulating feature directions [3, 10, 14, 31]. In these approaches, features in neural networks are represented as vectors in activation space, where each dimension can encode different aspects of the input data. This vectorial representation supports complex feature combinations, essential for the network’s ability to generalize from seen to unseen data [4]. The techniques allow the creation of attribution graphs that accurately map the dependencies among features [3, 8, 10].
Yet, the circuits extracted from neural networks using methods such as sparse autoencoders (SAEs) [8] and transcoders [10] are best understood as targeted approximations of the original model, rather than exact mappings of the underlying computational processes. Due to their nature, they require new trained components which inherently simplify or omit certain aspects of the original model’s complexity. This includes focusing on a subset of features, possibly ignoring less interpretable yet significant elements, and applying transformations like gating or thresholding that modify the original computational pathways. Furthermore, these methods often involve training new versions of network components—like using transcoders to replace dense layers with sparser ones. While this helps in understanding how inputs are processed, it is fundamentally a reconfiguration rather than a direct observation of the original network’s internal representation. In addition, recent work has raised doubts concerning the generalization ability of such approaches [19]. Consequently, while these extracted circuits are highly valuable for gaining insights into model behavior and assisting in debugging or educational purposes, they do not fully capture all the intricate details and interactions of the original model. In contrast, the idea behind our combinatorial interpretabiliy approach, though at a much earlier phase than the study of SAEs [8] and transcoders [10], is to directly interpret the computation from analyzing the structure of the model itself.
@@ 6 Modeling Computation and Combinatorial Disentanglement
Figure 20: The feature circuit coding-decoding view of computation. Computation flows from right to left. The middle sequence (b) is the actual computation through the neural network weight matrices. The top sequence (a) is the abstract computation mapping features to features in each layer being represented by the neural network. The lower sequence (c) is the feature coding view that bridges the two by thinking of the the weight matrix as implementing a decoding of the prior layer’s features to serve as monosemantic inputs to a coding-and-computation of the features of the new one. Interestingly, while $W_{i}=C_{i}D_{i}$ is a superposed $n_{i}\times n_{i-1}$ representation of the $m_{i}\times m_{i-1}$ $i$-th feature circuit layer, its also the case that if we put aside the $i$-th layer ReLU and bias calculations, then $D_{i+1}C_{i}$ is an $m_{i}\times m_{i-1}$ representation of this layer. Further explanations appear in the text body.
Our approach so far has been to show the combinatorial interpretation of a simple 2-layer network. In the real world, even multi-layered perceptrons (MLPs) are more complex, typically with an initial embedding [13, 29] and multiple layers. In this and the next sections we will provide a model and suggested path of how the combinatorial approach can be used to disentangle an MLP to expose its full feature circuit. We start with the model, show our initial progress in disentangling a network by showing how the initial embedding can be disentangled to expose the feature channel coding in the first layer, and then propose a technique for disentanglement that could work in its generality. Call this the cascading feature disentanglement technique.
@@ 6.1 The Feature Circuit View of Neural Computation
Our modeling follows the intuitive ideas in [1, 12, 39]. We capture neural network computation as a set of output features $F=\{f_{1},f_{2},...,f_{m}\}$, where each is the result of applying a computation to a set of input features [12]. We assume that $F$ is computed by a neural network $N(F)$. The feature circuit view of neural computation attempts to separate the computation from the chosen medium in which it is performed. To do so, we define a feature circuit as a sequence of computations $F_{1},F_{2},\ldots,F_{d}$.
An example feature circuit is depicted in the top part (a) of Figure 20: it is an abstract multi-layered computation that receives an input vector of $m_{0}$ inputs, and computes a cascading sequence of computations, where in each layer the output of a feature $f^{\prime}_{i}\in\{f^{\prime}_{1},f’_{2},...,f’_{m_{i}}\}$ is the result of applying $f^{\prime}_{i}$ to a subset of the input features $f_{1},f_{2},...,f_{m_{i-1}}$. In our simple 2-layer network examples in this paper, a fully connected layer followed by a single neuron, the feature circuit for a DNF of 2-AND features would look like the one depicted in Figure 21. Note that in that figure, the rectangular portion of the matrix depicts the dependencies (using filled in rectangles), with the actual computation being performed by the formulas to the left of the rectangle.
Figure 21: Two layer feature circuit (matching the top of Figure 20) of the DNF $(x_{3}\land x_{12})\lor\left(x_{2}\land x_{8}\right)\lor\left(x_{5}\land x_{11}\right)\lor\left(x_{0}\land x_{9}\right)\lor\left(x_{1}\land x_{6}\right)\lor\left(x_{10}\land x_{15}\right)\lor\left(x_{4}\land x_{13}\right)\lor\left(x_{7}\land x_{14}\right)$. Layer 1 defines the computation of the 2-AND features and Layer 2 os the single OR feature over the results of the outputs from Layer 1. Together they define the abstract computation that the network in Figure 22 computes.
This abstract computation has a concrete manifestation as a multi-layered perceptron, depicted in the middle part (b) of Figure 20, where the inputs are first encoded via an embedding $C_{0}$ to the first layer of neurons and then each layer $i$ is an $n_{i}\times n_{i-1}$ matrix $W_{i}$ defining the computation of $n_{i}$ neurons with a non-linearity (here ReLU) and a bias capturing the corresponding feature circuit computation for that layer. The network can have any number of layers $d$. For simplicity, we will refer to the non-linearity as a ReLU, although any other non-linearities (such as batch norm or max pooling) could be included in the model. The final output is a readout of the final layer via a softmax or similar function. This multi-layer perceptron, through the gradient descent based training process, is thus a manifestation of the original feature circuit up to a certain accuracy.
@@
As [1, 39] show, a neural network can use a polysemantic representation with many fewer neurons than features to compute across multi-level neural networks. Part of our observations in this work suggest that polysemanticity can arise in training even when the network does not need to be in superposition, that is, $n_{i}\geq m_{i}$. In some neural networks, such as large language models or convolutional neural networks, the input to the neural network is provided in superposition. In these cases, the input is encoded into superposition via a first layer that translates text (via tokens) into an embedding space or downsamples an image.
@@ 6.2 The Disentangled Feature Circuit View
A key element of our combinatorial approach to interpreting neural computation is to introduce a partly disentangled or “input disentangled” intermediate view of the concrete multi-layered perceptron. This is depicted in the lower part (c) of Figure 20, and is a step towards computing the actual feature circuit being represented. We view each layer’s weight matrix $W_{i}$ as the product of two matrices $C_{i}$ and $D_{i}$. $D_{i}$ is of dimension $m_{i-1}\times n_{i-1}$ and decodes the output features from the prior neuron layer of $n_{i-1}$ neurons into a monosemantic representation of those $m_{i-1}$ features. $C_{i}$ of size $n_{i}\times m_{i-1}$ in turn computes the $m_{i}$ different new features output by the current layer over those $m_{i-1}$ input features, and then projects their representation back into a coded representation on size $n_{i}$ for the next layer. The whole network is thus a sequence $C_{d}D_{d}\ldots C_{1}D_{1}$ of such decodings, opening the coded representation up to compute in feature space, and then coding to go back to the neuronal representation, all this with ReLU and bias computations between them and a softmax at the end. Each $W_{i}=C_{i}D_{i}$ so we only see an $n_{i}\times n_{i-1}$ matrix computation in the concrete neural network view: the actual feature circuit is hidden in $W_{i}$. We named the disentangled view matrices $D_{i}$ and $C_{i}$ also because $D_{i}$ is effectively decompressing the polysemantic representation across $n_{i-1}$ neurons into a typically larger set of $m_{i-1}$ monosemantic features, and in turn $C_{i}$ is compressing them back into a representation using $n_{i}$ polysemantic neurons.
Interestingly, if one ignores bias and ReLu, then notice that the the feature circuit for any layer $i$ could be given by $D_{i}*C_{i-1}$, a matrix of size $m_{i}\times m_{i-1}$ in which the codes of length $n_{i}$ are matched between $D_{i}$ and $C_{i-1}$ yielding the layer $i$ feature circuit matrix. So for example, as depicted in Figure 20, the Layer 1 feature circuit would be given by multiplying $D_{2}$ by $C_{1}$ (but of course this matrix is only an approximation because of the bias and ReLU, whereas $C_{1}D_{1}$ is an accurate representation of the neural network matrix $W_{1}$).
@@ 6.3 How the Feature Channels Code Computation
Let us spend a moment being more precise about how coding defines computation. As we noted, this is the real differentiator of feature channel coding from prior approaches. We assume that $f_{0}$, a feature to be computed, is a function of a small set of $k$ features $F=\{f^{\prime}_{1}\ldots f^{\prime}_{k}\}$ from the prior layer. The features $F$ are coded by the non-negative outputs of the prior layer, and $f_{0}=f_{0}(F)$ will be encoded by the non-zero outputs of the computation of the current layer, all other entries being 0. To perform this computation, the weight matrix $W_{i}$ maps each of the codes in $F$ to the code for $f$. When a feature at the prior layer is used for multiple features at the current layer, it gets mapped to each of the codes for the features it is used in. Thus, prior to the non-linearity for the activation at that layer, each of the inputs to $f_{0}(F)$ have been mapped to the same set of neurons at the layer representing $f_{0}$. This allows the network to use the code for $f_{0}$ as a computational channel for computing $f_{0}$. Unfortunately, this entangled mapping in $W_{i}$ is hiding the actual mapping $f_{0}(F)$. However, we know that the outputs of $D_{i}$ are the monosemantic (uncoded) input features in $F$, and the matrix $C_{i}$ defines the coding of $f_{0}(F)$ and defines the computation of $f_{0}(F)$ individually on each neuron of the code. Note that the channel’s code is somewhat orthogonal to the computation being performed - the goal of channel coding is to bring together the inputs that need to be combined to compute a function. Once they are on the same channel, a number of different functions can be performed on those incoming features. Thus, uncovering the codes for a given neural network would allow us to describe its circuit.
6.4 Decoding the Circuit
At the core of how we believe the combinatorial approach could work is the cascading disentanglement technique. We hypothesize that given a multi-layered neural network, we can decode the matrices $W_{i}$ to reveal the coding in each $C_{i}$ through a cascading sequence of decodings starting from the first layer and proceeding to the last, allowing us to reconstruct the complete description of the feature circuit.
Cascading Disentanglement for a MLP:
@@ 6.4 Decoding the Circuit
    Initially: One can decode $C_{1}$ as $C_{1}\approx W_{1}*C_{0}$.
@@
    For all $i>1$, to decode $C_{i}$, interpret the feature channel codes of the decoded $C_{i-1}$ to derive the Layer $i-1$ feature circuit, and use its outputs to construct $D_{i}^{-1}$. Then $C_{i}\approx W_{i}D_{i\,}^{-1}$. Repeat until you have decoded $C_{d}$.
Let us explain the Cascading Disentanglement technique in more detail. Lets start with the initial Step 1. As depicted in Figure 20, notice that $C_{1}\approx W_{1}*C_{0}$ because $D_{1}\approx C_{0}^{-1}$:  [^3] That is, $D_{1}$ is approximately the left inverse of $C_{0}$. We use the general form of matrix inverse here [40] that does not require symmetry. there is no bias or ReLU between the application of $C_{0}$ and $W_{1}$, and $C_{0}$ is an $n_{0}\times m_{0}$ matrix while $D_{1}$ is an $m_{0}\times n_{0}$ matrix.

### v2

@@ 1 Introduction
An exciting recent body of research in the interpretability of neural networks has been focused on viewing their computation as collections of independent features whose directions in high dimensional activation space can be analyzed. This has led, using novel approaches such as the use of sparse autoencoders [29], to the discovery of polysemanticity, the superposition hypothesis [12], and a body of empirical evidence of the correlation between features and directionality of neuronal activation space vectors [2, 3, 5, 7, 10, 11, 13, 29, 30]. Those techniques (surveyed in detail in the related work section) have provided ingenious ways of capturing the representation of features and some information on their interdependencies, but so far have not elucidated the full dependency graph, the underlying function being computed nor the mechanism for computing it.
@@ 2 Background and Related Work
Recent advances in the use of autoencoders for network interpretability have led to enhanced methods for isolating and manipulating feature directions [3, 10, 13, 29]. In these approaches, features in neural networks are represented as vectors in activation space, where each dimension can encode different aspects of the input data. This vectorial representation supports complex feature combinations, essential for the network’s ability to generalize from seen to unseen data [4]. The techniques allow the creation of attribution graphs that map the dependencies among features for the computation of a specific input [3, 8, 10].
Yet, the circuits extracted from neural networks using methods such as sparse autoencoders (SAEs) [8] and transcoders [10] are best described as targeted approximations of the original model, rather than elucidating the underlying computational processes. Specifically, the resulting circuits capture which features influence subsequent features for a specific input, but they do not capture (a) the overall feature dependence (across all inputs), (b) the function that is being computed to realize that dependence, and (c) the underlying mechanism for computing that function.
Due to their nature, SAE-like techniques require new trained components which inherently simplify or omit certain aspects of the original model’s complexity. This includes focusing on a subset of features, possibly ignoring less interpretable yet significant elements, and applying transformations like gating or thresholding that modify the original computational pathways. Furthermore, these methods often involve training new versions of network components—like using transcoders to replace dense layers with sparser ones. While this helps in understanding how inputs are processed, it is fundamentally a reconfiguration rather than a direct observation of the original network’s internal representation. In addition, recent work has raised doubts concerning the generalization ability of such approaches [18]. Consequently, while these extracted circuits are highly valuable for gaining insights into model behavior and assisting in debugging or educational purposes, they do not fully capture all the intricate details and interactions of the original model. In contrast, the idea behind our combinatorial interpretabiliy approach, though at a much earlier phase than the study of SAEs [8] and transcoders [10], is to directly interpret the computation from analyzing the structure of the model itself.
@@ 6 Modeling Computation and Combinatorial Disentanglement
Figure 20: The feature circuit coding-decoding view of computation. Computation flows from right to left. The middle sequence (b) is the actual computation through the neural network. The top sequence (a) is the abstract computation mapping features to features in each layer being represented by the neural network. The lower sequence (c) is the feature coding view that bridges the two by thinking of the the weight matrix as implementing a decoding of the prior layer’s features to serve as monosemantic inputs to a coding-and-computation of the features of the new one. $W_{i}=C_{i}D_{i}$ is a superposed $n_{i}\times n_{i-1}$ representation of the $m_{i}\times m_{i-1}$ $i$-th feature circuit layer.
Thus far we have illustrated the combinatorial view of a two‑layer network. Real‐world models—even plain multilayer perceptrons (MLPs)—are usually richer: they begin with an embedding layer and then stack several nonlinear layers. This section introduces a conceptual model for understanding MLP computation based on underlying “feature circuits.” We then propose a novel perspective viewing MLP layers as implicitly performing a decode-compute-encode cycle, aiming to disentangle the network and reveal this circuit. Building on this, we outline a potential strategy, termed cascading feature disentanglement, for systematically uncovering the feature circuit layer by layer. We present initial steps demonstrating how the first layer of computation can be revealed via disentanglement from an input embedding and also propose a general approach for deeper layers.
@@ 6.1 The Feature Circuit View of Neural Computation
Our modeling follows the intuitive ideas in [1, 12, 37]. We aim to capture the fundamental computational structure implemented by an MLP. We model the computation at each layer $i$ (for $i=1,\ldots,d$) as producing a set of abstract output features $F_{i}=\{f_{i,1},f_{i,2},\ldots,f_{i,m_{i}}\}$. Each feature $f_{i,j}\in F_{i}$ is computed by applying a specific function to a subset of the features $F_{i-1}=\{f_{i-1,1},\ldots,f_{i-1,m_{i-1}}\}$ from the preceding layer (where $F_{0}$ represents the initial input features, $m_{0}$ in number). The complete feature circuit of the network is the sequence of these computations across all layers, $F_{1},F_{2},\ldots,F_{d}$.
Figure 20(a) provides a visualization of such a multi-layer feature circuit. It takes $m_{0}$ input features and computes features layer by layer, where the computation of a feature $f_{i,j}$ depends only on a subset of features from $F_{i-1}$. For instance, in our earlier simple 2-layer network example (a fully connected layer followed by a single neuron computing a DNF of 2-ANDs), the feature circuit might resemble Figure 21. Note that in that figure, the rectangular portion of the matrix depicts the dependencies (using filled in rectangles), and the formulas to the left of the rectangle depict the actual functions being computed.
Figure 21: Two layer feature circuit (as an example of the top row of Figure 20) of the DNF $(x_{3}\land x_{12})\lor\left(x_{2}\land x_{8}\right)\lor\left(x_{5}\land x_{11}\right)\lor\left(x_{0}\land x_{9}\right)\lor\left(x_{1}\land x_{6}\right)\lor\left(x_{10}\land x_{15}\right)\lor\left(x_{4}\land x_{13}\right)\lor\left(x_{7}\land x_{14}\right)$. Layer 1 defines the computation of the pairwise AND features and Layer 2 is the single OR feature over the results of the outputs from Layer 1. Together they define the abstract computation that the network in Figure 22 computes.
This abstract feature circuit finds a concrete implementation in a standard MLP, depicted in Figure 20(b). The $m_{0}$ input features are first mapped by an embedding matrix $C_{0}$ (size $n_{0}\times m_{0}$) to an $n_{0}$-dimensional vector. Subsequent layers $i=1,\ldots,d$ consist of a weight matrix $W_{i}$ (size $n_{i}\times n_{i-1}$), a bias vector $b_{i}$, and a non-linearity (e.g., ReLU), transforming the $n_{i-1}$ neuron activations from layer $i-1$ into $n_{i}$ neuron activations for layer $i$. (While we use ReLU for illustration, other operations like batch normalization could be used instead). A final readout function (e.g., softmax) produces the network’s output. Through training (e.g., via gradient descent), this MLP learns to approximate the underlying feature circuit computation.
@@
As [1, 37] show, a neural network can use a polysemantic representation with many fewer neurons than features to compute across multi-level neural networks. Part of our observations in this work suggest that polysemanticity can arise in training even when the network does not need to be in superposition, that is, $n_{i}\geq m_{i}$ (similar to results shown in [22]). In some neural networks, such as large language models or convolutional neural networks, the input to the neural network is provided in superposition. In these cases, the input is encoded into superposition via a first layer that translates text (via tokens) or an image (via downsampling) into an embedding space.
@@ 6.2 The Disentangled Feature Circuit View
Our core proposal for interpreting MLPs involves viewing each layer’s weight matrix $W_{i}$ as implicitly factoring into two conceptual matrices, $C_{i}$ and $D_{i}$, revealing an intermediate “disentangled” representation, shown in Figure 20(c). This perspective aims to bridge the gap between the concrete MLP and the abstract feature circuit.
Specifically, we hypothesize that the computation within layer $i$ can be understood as:
  1. 1.
    Decompression ($D_{i}$): An implicit matrix $D_{i}$ (size $m_{i-1}\times n_{i-1}$) acts on the $n_{i-1}$ neuron activations from layer $i-1$. Its role is to transform the $m_{i-1}$ polysemantic features computed by the previous layer ($F_{i-1}$) into approximations of their monosemantic representations.
  2. 2.
    Computation and Compression ($C_{i}$): An implicit matrix $C_{i}$ (size $n_{i}\times m_{i-1}$), combined with the bias $b_{i}$ and non-linearity $\sigma_{i}$ (e.g., ReLU), acts on the monosemantic representations. It performs two roles simultaneously, using feature channel coding:
    • It computes the $m_{i}$ new features for the current layer ($F_{i}$) based on the $m_{i-1}$ input features.
    • It compresses these $m_{i}$ computed features back into the $n_{i}$-dimensional polysemantic represenations for the next layer.
Thus, the effective transformation by the weight matrix $W_{i}$ in the standard MLP (Figure 20(b)) is conceptually equivalent to the combined operation $W_{i}\approx C_{i}D_{i}$ in our disentangled view (Figure 20(c)). The entire network can then be viewed as a sequence
$C_{d}D_{d}\sigma_{d-1}\ldots\sigma_{1}C_{1}D_{1}(C_{0}\cdot\text{input}),$
where $\sigma_{i}$ represents the bias and non-linearity applied after the $C_{i}D_{i}$ (or equivalently, $W_{i}$) multiplication. The actual MLP uses the compact $W_{i}=C_{i}D_{i}$ form, hiding the intermediate dimension of size equal to the number of features at layer $i$. The names $D_{i}$ (Decompress) and $C_{i}$ (Compress) reflect their roles in moving between the compact polysemantic representation and the potentially larger monosemantic representation.
This perspective is similar to concepts like sparse autoencoders and transcoders, where encode/decode operations learn similar transformations. Specifically, the $D_{i}$ matrix resembles an encoder and the $C_{i}$ matrix a decoder of a transcoder. However, a key distinction is that $C_{i}$ and $D_{i}$ are posited as intrinsic components of the existing trained weight matrix $W_{i}$, not externally added modules learned via a separate objective like in standard autoencoders or transcoders. Our goal is to recover this inherent $C_{i}D_{i}$ structure. Consequently, the placement of non-linearities and biases in our model follows the original network architecture, differing from typical autoencoder/transcoder setups.
Interestingly, if one ignores bias and ReLu, then notice that the feature circuit for any layer $i$ could be given by $D_{i}*C_{i-1}$, a matrix of size $m_{i}\times m_{i-1}$ in which the codes of length $n_{i}$ are matched between $D_{i}$ and $C_{i-1}$ yielding the layer $i$ feature circuit matrix. So for example, as depicted in Figure 20, the Layer 1 feature circuit would be given by multiplying $D_{2}$ by $C_{1}$. This matrix is only an approximation because of the bias and ReLU, whereas $C_{1}D_{1}$ is an accurate representation of the neural network matrix $W_{1}$.
@@ 6.3 How the Feature Channels Code Computation
Let us spend a moment revisiting how feature channel coding defines computation in this view of the neural network. We assume that $f_{i,j}$, a feature to be computed at layer $i$, is a function of a small set of $k$ features $G=\{f_{i-1,1}\ldots f_{i-1,k}\}$ from layer $i-1$. The features in $G$ of the feature circuit view are each coded (n the neural network view) by a set of neurons from layer $i-1$, and $f_{i,j}$ is coded by a set of neurons from layer $i$. To perform the computation for $f_{i,j}$, the weight matrix $W_{i}$ maps each of the codes for features in $G$ to the code for $f_{i,j}$. When a feature at layer $i-1$ is used for multiple features at layer $i$, it gets mapped to each of the codes for the features it is used in. Thus, prior to the non-linearity for the activation at layer $i$, each of the inputs to $f_{i,j}$ have been mapped to the same set of neurons at layer $i$. This allows the network to use the code for $f_{i,j}$ as a computational channel for computing $f_{0}$.
The entangled matrix $W_{i}$ performs this mapping for the features used by $f_{i,j}$. The disentangled view makes this clear: $D_{i}$ first maps those features to their monosemantic representation. $C_{i}$ then brings them all together to the code for $f_{i,j}$. In a funny way, in terms of feature channel coding, $D_{i}$ is a decoding matrix, moving codes into features, and $C_{i}$ is a coding matrix, encoding them back into new collections of features, exactly the opposite of the naming convention in the autoencoding/transcoding literature [29, 10]. The matrix $C_{i}$, combined with the bias and the non-linearity, performs the computation required for $f_{i,j}$. In the examples provided in earlier sections, the inputs were already in their monosemantic representation, so there was no need for a $D$ matrix; the layer one weight matrix serves directly as the $C$ matrix as well. Note that the channel’s code is somewhat independent of the computation being performed - the goal of channel coding is to bring together the necessary inputs onto the same computational channel. Once they are on the same channel, a number of different functions can be performed on those incoming features. Thus, uncovering the codes for a given neural network would allow us to describe its feature circuit.
6.4 Cascading Feature Disentanglement
We propose the cascading feature disentanglement technique as a potential avenue for recovering the feature circuit from a trained MLP. The core hypothesis is that we can iteratively uncover the matrices $C_{i}$ and $D_{i}$ layer by layer.
Cascading Disentanglement Technique: For a $d$-layer MLP with weights $W_{1},\ldots,W_{d}$ and input embedding $C_{0}$:
@@ 6.4 Cascading Feature Disentanglement
    Initialization (Layer 1): Estimate $C_{1}$. Since there is typically no non-linearity between the input embedding $C_{0}$ (size $n_{0}\times m_{0}$) and the first weight matrix $W_{1}$ (size $n_{1}\times n_{0}$), the initial disentanglement is simply $C_{1}\approx W_{1}C_{0}$. The role of $D_{1}$ (size $m_{0}\times n_{0}$) is to invert the input embedding, i.e., $C_{0}=D_{1}^{*}$ (where $D_{1}^{*}$ denotes an approximate right inverse of $D_{1}$, as $D_{1}$ is generally non-square [38]). If this holds, then since $W_{1}\approx C_{1}D_{1}$, we see that $W_{1}C_{0}\approx C_{1}D_{1}C_{0}=C_{1}D_{1}D_{1}^{*}\approx C_{1}$.
@@
    Iteration (Layer $i>1$): Assume we have successfully estimated $C_{i-1}$. Our goal is to estimate $C_{i}$ from $W_{i}\approx C_{i}D_{i}$. Unlike Step 1, we cannot simply assume $D_{i}\approx C_{i-1}^{*}$ because a non-linearity and bias ($\sigma_{i-1}$) were applied after the computation involving $C_{i-1}$. The central hypothesis for this step is that by analyzing the estimated $C_{i-1}$ (size $n_{i-1}\times m_{i-2}$), we can potentially deduce the $m_{i-1}$ features computed by layer $i-1$ and how they are represented by the $n_{i-1}$ neuron activations (after bias and ReLU). This understanding of the layer $i-1$ output features could allow us to construct an estimate for $D_{i}$ (and hence its approximate inverse $D_{i}^{*}$). If we can find $D_{i}^{*}$, we can then estimate $C_{i}\approx W_{i}D_{i}^{*}$. Repeating this process for $i=2,\ldots,d$ would, in principle, decode the entire network’s feature circuit via $C_{1},\ldots,C_{d}$.
Developing the precise methods to infer the layer $i-1$ feature vectors from $C_{i-1}$ and construct the corresponding $D_{i}$ (or $D_{i}^{*}$) remains a significant challenge and is a key direction for future research. This cascading approach, however, provides a conceptual roadmap for systematically disentangling deep MLPs. [^3] Note that we are proposing an algorithmic framework without having the solution algorithm, apart from some initial heuristic attempts. Skeptics might criticize us for this. We note that this has been done before, for example, in the context of public key cryptography [9] where Diffie and Hellman proposed a public key cryptography algorithm without the RSA algorithm [31] to implement it. We also point out that an alternative approach would be to use a sparse autoencoder to determine the feature encoding being input to layer $i$, and use that to construct the pseudo-inverse $D_{i}^{*}$. In fact, the decode matrix of the sparse autoencoder may itself serve as the pseudo-inverse $D_{i}^{*}$, since it essentially performs the same function for layer $i$ as an initial embedding matrix does for layer 1.
6.5 Cascading Feature Disentanglement with an Embedding Layer
We demonstrate that that Cascading Disentanglement works for a small example network where we add an embedding layer to the two layer networks we have described earlier in the paper. Specifically, Figure 22 depicts three example networks trained via gradient descent, where three different types of embeddings are added to the feature circuit depicted in Figure 21. Each row of matrices is a different type of embedding, where the rightmost column depicts the embedding being used: in the middle row, the embedding $C_{0}$ is the identity function (i.e., no embedding), in the top row it is a Hadamard matrix, and the bottom row it is a symmetric random embedding.
The middle column depicts the raw matrix $W_{1}$, i.e., prior to disentanglement. In the case of the identity embedding (middle row) feature channel coding is clear. In the case of the Hadamard and random embeddings, the $W_{1}$ matrices look random. The leftmost column depicts the result of disentanglement: the matrix obtained by the product $W_{1}C_{0}$. As described above, this gives us $C_{1}$ since $W_{1}C_{0}\approx C_{1}D_{1}C_{0}\approx C_{1}D_{1}D_{1}^{*}\approx C_{1}$. This reveals the feature channel encoding hidden within $W_{1}$ for both the Hadamard embedding and the random embedding. In other words, multiplying by $C_{0}$ undoes the implicit embedding that was learned by the network in $W_{1}$, and reveals the feature channel codes.